what is the slope line through the points (2,2) and (4,3)?
The answer is 1/2 but i don't understand how?

Answer by dabanfield(803)

what is the slope line through the points (2,2) and (4,3)?
The answer is 1/2 but i don't understand how?

The formula for the slope of the line between two points (x1,y1) and (x2,y2) is:
(y2-y1)/(x2-x1)
So in this case the slope is:
(3-2)/(4-2) = 1/2
Notice that if we reversed the points would also have:
(2-3)/(2-4) = -1/-2 = 1/2